The price of widgets is $20 and the quantity demanded at $20 is 100. The price of widgets decreases to $10 and the quantity demanded increases to 120. Calculate the price elasticity of demand, being sure to show enough work that it is clear to follow your process. Is this elastic or inelastic? What should be done, if anything, to price to increase revenues?
Answer #1

Price elasticity of demand = % change in the demand / % change in the price.

% change in the demand = (120 -100) / (120 +100)

= 20 / 220

= 0.09

% change in the price = (10-20) / (10+20)

= -10 / 30

= -0.33

Price elasticity = 0.09 / -0.33

= -0.24

As the value of elasticity is below 1 the demand is inelastic. to increase the revenue the seller need to increase the price.